The Basket Weavers Company has 100,000 units of semi-annual coupon, 20-year bonds outstanding that are currently selling at par value ($1000). The coupon rate of the bond is 7.47%. The company also has 1 million shares of 10.5 percent preferred stock outstanding and 5 million shares of common stock outstanding. The preferred stock has a par value of $100 and is selling for $60 per share. The common stock has a beta of 1.5 and is selling for $40 per share. The U.S. Treasury bill is yielding 3 percent and the return on the market portfolio is 12 percent. The corporate tax rate is 35 percent.

(a) Calculate the cost of equity of the Basket Weavers Company. (2 marks)

.

(b) Calculate the cost of preferred stock of the Basket Weavers Company. (2 marks)

.

(c) Calculate the after-tax cost of debt of the Basket Weavers Company. (2 marks)

.

(d) Calculate the weighted average cost of capital of the Basket Weavers Company. (5 marks)

.

(e) Should the weighted average cost of capital be used as a required return for an investment project which has a much higher risk than the overall risk of the company? Explain briefly.
Answer #1
a] Cost of preferred stock = 10.5/60 = 17.50%
b] After tax cost of debt = 7.47%*(1-35%) = 4.86%
c] Cost of equity per CAPM = 3%+1.5*(12%-3%) = 16.50%
d] The WACC using market value weights is calculated in the table below:
Source of capital Market Value [$ Million] Weight Before Tax Cost Weighted Cost
Debt [100000*1000] $              100.00 27.78% 4.86% 1.35%
Preferred stock [1000000*60] $                60.00 16.67% 17.50% 2.92%
Common stock [5000000*40] $              200.00 55.56% 16.50% 9.17%
Total $              360.00 13.43%
WACC 13.43%
e] The WACC should be used as the required return only for those projects that have the same risk as the existing
business of the firm. If not so, the WACC has to be adjusted to represent the true risk of the proposed project.
Here, the risk of the proposed is higher and hence the WACC has to be adjusted upwards to represent the
higher risk of the project.